OpenStudy (anonymous):

Is possible to solve the integral of 1/(a/x -b) . If yes, which is the solution?

OpenStudy (anonymous):

\[\int \frac 1{\frac ax - b}\] right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you can simplify it to \[\int \frac 1{\frac{a-bx}x}\] and then \[\int \frac x{a-bx}\] does that help you see the integration now?

OpenStudy (anonymous):

ok, do you have the solution for this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Would you mind posting it?

OpenStudy (anonymous):

Substitution: \(u=a-bx~\Rightarrow~du=-dx\). Also, \(u=a-bx~\iff~x=\dfrac{a-u}{b}\). So you have \[\begin{align*}\int\frac{x}{a-bx}~dx&=-\int\frac{\frac{a-u}{b}}{u}~du\\ &=\int\frac{u-a}{bu}~du\\ &=\int\left(\frac{1}{b}-\frac{a}{bu}\right)~du \end{align*}\]